# Compressed data structures for Heegaard splittings

### v1

@@ Abstract.
Heegaard splittings provide a natural representation of closed 3-manifolds by gluing handlebodies along a common surface. These splittings can be equivalently given by two finite sets of meridians lying in the surface, which define a Heegaard diagram. We present a data structure to effectively represent Heegaard diagrams as normal curves with respect to triangulations of a surface of complexity measured by the space required to express the normal coordinates’ vectors in binary. This structure can be significantly more compressed than triangulations of 3-manifolds, given exponential gains for some families. Even with this succinct definition of complexity, we establish polynomial time algorithms for comparing and manipulating diagrams, performing stabilizations, detecting trivial stabilizations and reductions, and computing topological invariants of the underlying manifolds, such as their fundamental and first homology groups. We also contrast early implementations of our techniques with standard software programs for 3-manifolds, achieving better precision and faster algorithms for the average cases and exponential gains in speed for some particular presentations of the inputs.
@@ 1. Introduction
[…]
@@
Heegaard splittings are one of these alternatives. They depict a closed 3-manifold $M$ as the result of gluing two better-behaved pieces called handlebodies along a common boundary. Heegaard splittings can also be combinatorially described by Heegaard diagrams, two sets of curves in a surface whose intersection pattern conveys a lot of topological information about $M$. We suggest here a new data structure to represent Heegaard diagrams based on normal coordinates, of a size as large as the binary of the number of intersection points between the two sets of curves. The gain in measuring the complexity of 3-manifolds through this notion is not to be overlooked: we present in Remark 5.1 families of manifolds whose complexities of diagrams grow linearly on the size of the input, but whose (direct) triangulations use exponentially many tetrahedra. What is more important, however, is that due to the proliferation of efficient operations on normal coordinates in the last two decades [18, 34, 48, 49, 55], we are able to establish polynomial time algorithms even in this compressed notion of complexity, leading to basic manipulations of Heegaard splittings and 3-manifolds. We hope that these techniques are useful for the study and classification of topological spaces that are beyond the reach of efficiency when represented by other methods.
@@
This paper starts with a background section on both low-dimensional and computational topology (Section 2), where we review some of the concepts and results that we will need in later sections. To keep this paper short, we refer the reader to [52] and [54] for basic ideas of geometric topology, such as homeomorphisms, isotopies, gluing maps, and manifolds. Our contribution starts in Section 3, where we formally present our data structure for Heegaard splittings. Section 4 establishes some basic algorithms in this structure. We then compare early implementations [^1] The code is publicly available at https://github.com/HLovisiEnnes/FHDpy. Details on the implementation are also outlined in Appendix B. of these algorithms with triangulation-dependent methods in Section 5. We also present appendices with details on the implementation and extensions of the techniques described in the main text.
@@
Before we close this introduction, we call the reader’s attention to two subtleties of the current work. First, although most of our tools are applicable to generalized Heegaard splittings defined for orientable compact 3-manifolds, potentially with boundary [31, 52], we will solely focus on closed 3-manifolds. Therefore, all 3-manifolds here will be assumed orientable and closed, unless explicitly stated otherwise. Second, we assume throughout a unit-cost RAM model, with some bit-size $W$ [20]. This means that the four integer operations and integer comparisons can be executed in constant time, provided that the numbers involved have binary representations of size at most $W$.
@@ 2.1. Background on low-dimensional topology
By surface, we understand a compact and connected smooth 2-manifold $\Sigma$ with or without boundary. All surfaces will be assumed orientable and oriented. If $\Sigma$ is closed, it is uniquely given, up to homeomorphism, by an integer $g$ called the surface’s genus [19] and we denote this surface by $\Sigma_{g}$. A (simple closed) curve and an arc are images of proper embeddings of $S^{1}$ and $[0,1]$ in $\Sigma$, respectively. A multicurve is a finite collection of disjoint curves. Given a multicurve $\gamma$, we denote its number of connected components by #$\#\gamma$ and, with some abuse of notation, write $c\in\gamma$ if $c$ is a component of $\gamma$. Whenever possible, we reserve Greek letters for multicurves and Latin letters for curves and arcs.
We will most often consider the isotopy class of a (multi)curve $\gamma$ in $\Sigma$, $[\gamma]$. A class $[\gamma]$ is called essential if no component is isotopic to point. From now on, we will not differentiate $\gamma$ from its isotopy class $[\gamma]$ in the notation if no confusion arises.
Let $\gamma$ and $\delta$ be two oriented (multi)curves in $\Sigma$, isotoped to intersect transversely. For each intersection point, we let its index be $+1$ if the orientation of the curves agrees with the orientation of $\Sigma$ as in the right-hand rule, otherwise we set it to $-1$. The algebraic intersection number of $\gamma$ and $\delta$, $\hat{i}(\gamma,\delta)$, is the sum of the indices of all their transverse intersections; the algebraic intersection number is well defined up to isotopy [19]. Two (multi)curves are said to be in efficient position if they are representative of their respective isotopic classes such that the total number of intersection points between them is minimal. Equivalently, they are in efficient position if they do not have segments that cobound disks, known as bigons, in the surface [19].
The mapping class group of the surface $\Sigma_{g}$, $\operatorname{Mod}(\Sigma_{g})$, is the group of orientation-preserving $\Sigma_{g}\to\Sigma_{g}$ homeomorphisms up to isotopy. For each $g$, $\operatorname{Mod}(\Sigma_{g})$ is finitely generated as a group by an alphabet $\mathcal{L}$, consisting of Dehn twists about $13g-$ 1 (isotopy classes of) canonically defined essential curves $s_{1},\dots,s_{3g-1}$ [36]. Formally, a Dehn twist $\tau_{s}:\Sigma_{g}\to\Sigma_{g}$ about $s$ is given by cutting out a cylindrical regular neighborhood around the curve, giving a $2\pi$ left twist to one of the cylinder’s boundaries and gluing the cylinder back into the surface. We call the curves $s_{1},\dots,s_{3g-1}$ the surface’s Lickorish generators. Hopefully without confusion, we also call the Dehn twists in the alphabet =$\mathcal{L}=\{\tau_{s_{1}},\dots,\tau_{s_{3g-1}}\}$ Lickorish generators.
@@
[…]
where we attach 2-handles along the curves $\gamma$ in $\Sigma\times\{0\}$ and fill any spherical boundary component with 3-handles. Note that $\partial\mathcal{H}_{\gamma}=\Sigma_{g}$ and that the handlebody $\mathcal{H}_{\gamma}$ is homeomorphic to the connected sum of $g$ solid tori. The components of $\gamma$ are meridians of $\mathcal{H}_{\gamma}$ in the sense that they bound disks in the handlebody’s interior. Two systems $\gamma$ and $\gamma^{\prime}$ on the same surface $\Sigma_{g}$ will be called equivalent if $\gamma^{\prime}$ is a set of meridians in $\mathcal{H}_{\gamma}$ (which implies that $\gamma$ is a set of meridians in $\mathcal{H}_{\gamma^{\prime}}$).
@@
In this paper, we restrict the analysis to minimal systems $\gamma$, that is, we assume #$\#\gamma=g$. In recent decades, it became common to extend systems to any multicurve in the surface that defines a handlebody [25, 56]. In fact, the data structure described in Section 3 can be generalized to this notion of system and there are algorithms to translate from and to the special case of cardinality $g$. For completeness, we describe these extensions in the Appendix A.
Let $c_{1},c_{2}$ be two fixed components of a system $\gamma$ and $p$ an arc from a puncture corresponding to $c_{1}$ to a puncture corresponding to $c_{2}$ in $\Sigma_{g}\backslash\gamma$. We define the band sum of $c_{1},c_{2}$ and $p$ as the curve $c$ in $\Sigma_{g}\backslash\gamma$ indicated by the surgery in Figure 1. One can show that $c$ is essential and bounds a disk in $\mathcal{H}_{\gamma}$ [16, 31], implying that $\gamma^{\prime}=\gamma-\{c_{1}\}\cup\{c\}$ is a system equivalent to $\gamma$. We call this whole operation a disk slide. With finitely many applications of disk slides and isotopies, a system $\beta$ can be transformed into any equivalent system $\beta^{\prime}$ [31].
@@
A genus $g$ Heegaard splitting is the closed 3-manifold resulting from a gluing of two handlebodies with boundaries homeomorphic to $\Sigma_{g}$ through a gluing map $f:\Sigma_{g}\to\Sigma_{g}$. In fact, one might assume that $f$ is given by a $\phi\in\operatorname{Mod}(\Sigma_{g})$ composed with some canonical orientation-reversing homeomoprhism. This means that a Heegaard splitting can be given by a pair $(\Sigma_{g},\phi)$ where $\phi$ is a word on Lickorish generators. We call this presentation of Heegaard splittings a Heegaard word. Every closed 3-manifold has a Heegaard splitting and can then be associated with some pair $(\Sigma_{g},\phi)$ [43, 47].
@@
Although naturally combinatorial, a Heegaard word $(\Sigma_{g},\phi)$ hardly conveys topological information about the closed 3-manifold $M$ that it represents. Nonetheless, if we fix a system of meridians $\alpha$ on one of the two handlebodies, $\mathcal{H}_{\alpha}$, then $\beta=\phi(\alpha)$ is a system of meridians of the other handlebody, $\mathcal{H}_{\beta}$. We call the tuple $(\Sigma_{g},\alpha,\beta)$ a Heegaard diagram. We note that Heegaard diagrams are algorithmically independent from Heegaard words: any two systems $\alpha$ and $\beta$ will define the same splitting, provided that the curves are meridians in their respective handlebodies. Diagrams are therefore not unique for a given splitting. For example, if $\beta$ and $\beta^{\prime}$ are isotopic diagrams on $\Sigma_{g}$—that is, $\beta$ and $\beta^{\prime}$ are isotopic multicurves in $\Sigma_{g}$—then $(\Sigma_{g},\alpha,\beta)$ and $(\Sigma_{g},\alpha,\beta^{\prime})$ naturally define the same manifold. Similarly, if $\beta$ is equivalent to $\beta^{\prime}$, we say that the diagrams are equivalent and note that $(\Sigma_{g},\alpha,\beta)$ and $(\Sigma_{g},\alpha,\beta^{\prime})$ also define the same 3-manifold.
@@
Stabilizations provide of yet another method to generate alternative Heegaard splittings of a same 3-manifold. Formally, let $(\Sigma_{g_{1}},\alpha_{1},\beta_{1})$ be a Heegaard diagram of a 3-manifold $M_{1}$ and $(\Sigma_{g_{2}},\alpha_{2},\beta_{2})$ a diagram of $M_{2}$. The diagram $(\Sigma_{g_{1}+g_{2}},\alpha_{1}\cup\alpha_{2},\beta_{1}\cup\beta_{2})$ represents the connected sum $M_{1}\#M_{2}$, where $\Sigma_{g_{1}+g_{2}}$ is the connected sum of $\Sigma_{g_{1}}$ and $\Sigma_{g_{2}}$ taken along open disks disjoint from $\alpha_{1}$,$\alpha_{2}$, $\beta_{1}$, and $\beta_{2}$. In particular, noting that $S^{3}$ has a splitting with diagram $(\Sigma_{1}=T^{2},m,\ell)$, where $|m\cap\ell|=1$, and that $M\#S^{3}$ is homemorphic to $M$, the splitting $(\Sigma_{g+1},\beta\cup\{\ell\})$ represents the same manifold manifold as $(\Sigma_{g},\alpha,\beta)$. We say that the former is a stabilization of the latter. In the other direction, a Heegaard diagram $(\Sigma_{g},\alpha,\beta)$ represents a stabilization of a lower genus splitting if and only if there is a meridian $a$ in $\mathcal{H}_{\alpha}$ and a meridian $b$ in $\mathcal{H}_{\beta}$ such that $|a\cap b|=1$ [31]. In particular, we say that this pair of meridians is a stabilization pair and that the diagram is stabilized. For a Heegaard diagram $(\Sigma_{g},\alpha,\beta)$, we say that $a$ and $b$ form a trivial stabilization pair if $a$ is a component $\alpha$, $b$ is a component of $\beta$, they form a stabilization pair, and $b\cap a^{\prime}=\emptyset$ and $b^{\prime}\cap a=\emptyset$ for any other component $a^{\prime}\in\alpha$ and $b^{\prime}\in\beta$. Whenever $a$ and $b$ form a trivial stabilization, cutting the surface along $a$ induces a Heegaard diagram in the surface of genus one unit smaller. This process will be called destabilization.
@@
A splitting $(\Sigma_{g},\mathcal{H}_{\alpha},\mathcal{H}_{\beta})$ is called reducible if there is a curve $c$ in $\Sigma_{g}$ that is a meridian of both $\mathcal{H}_{\alpha}$ and $\mathcal{H}_{\beta}$. We call a diagram $(\Sigma_{g},\alpha,\beta)$ reducible if either (1) there are a component $a\in\alpha$ and a component $b\in\beta$ that are isotopic in $\Sigma_{g}$; or (2) if there is an essential separating curve $s$, disjoint from both $\alpha$ and $\beta$, and such that $\Sigma_{g}\backslash s$ defines a connected sum of two Heegaard splittings. The uniqueness of fundamental group decomposition [3] and [52][Theorem 6.3.5] imply that a reducible diagram represents a reducible splitting. A 3-manifold is reducible if it contains a sphere that does not bound a ball. If $(\Sigma_{g},\alpha,\beta)$ is stabilized, then either the splitting is reducible or a stabilization of a genus 1 splitting of $S^{3}$; moreover, a reducible splitting is stabilized or the splitting of a reducible manifold [31].
@@
Useful algebraic information on the 3-manifold can also be readily obtained from its Heegaard diagrams. Label the curves $a_{1},\dots,a_{g}\in\alpha$ and $b_{1},\dots,b_{g}\in\beta$ and give each of them an orientation. It follows from an application of the Van Kampen theorem [22, 26, 31] that we can compute a presentation $\langle x_{1},\dots,x_{g}|r_{1},\dots,r_{g}\rangle$ for the fundamental group $\pi_{1}(M)$ by traversing the curve $b_{i}$ along the assumed orientation and, whenever it intersects some $a_{j}$, we append $x_{j}$ to the relation $r_{i}$ if the index of the intersection is positive, otherwise we append $x_{j}^{-1}$ to it (if $b_{i}$ does not intersect $\alpha$, $r_{i}$ is just the trivial relation).
Now, recall that the first (integral) homology group of $M$, $H_{1}(M)$, is isomorphic to the abelianization of $\pi_{1}(M)$ [22]. We may therefore present $H_{1}(M)$ as
$\langle x_{1},\dots,x_{g}|r^{\prime}_{1},\dots,r^{\prime}_{g},[x_{i},x_{j}],1\leq i<j\leq g\rangle,$
@@
where each $r^{\prime}_{i}=x_{1}^{k_{i}^{1}}\dots x_{g}^{k_{i}^{g}}$ can be computed by commuting the generators $x_{j}$ in the relations $r_{i}$. Note that $k_{i}^{j}=\hat{i}(b_{i},a_{j})$ and we store this information in a $g\times g$ presentation matrix for $H_{1}(M)$, $K=[k_{i}^{j}]$, whose Smith normal form carries all the necessary information about this Abelian group. For example, the (first) Betti number of $M$ is given by the number of zero-rows. A 3-manifold with Betti number equal to zero is called a rational homology sphere, and it will be an integral homology sphere if the Smith form of the presentation matrix is the identity.

### v2

@@ Abstract.
Heegaard splittings provide a natural representation of closed 3-manifolds by gluing two handlebodies along a common surface. These splittings can be equivalently given by two finite sets of meridians lying on the surface, which define a Heegaard diagram. We present a data structure to effectively represent Heegaard diagrams as normal curves with respect to triangulations of a surface, where the complexity is measured by the space required to express the normal coordinates’ vectors in binary. This structure can be significantly more compact than triangulations of 3-manifolds, yielding exponential gains for certain families. Even with this succinct definition of complexity, we establish polynomial-time algorithms for comparing and manipulating diagrams, performing stabilizations, detecting trivial stabilizations and reductions, and computing topological invariants of the underlying manifolds, such as their fundamental and homology groups. We also contrast early implementations of our techniques with standard software programs for 3-manifolds, achieving faster algorithms for the average cases and exponential gains in speed for some particular presentations of the inputs.
@@ 1. Introduction
Since the early days of computational topology, 3-manifolds have attracted researchers’ keen interest [haken1962homoomorphieproblem, jaco1984algorithm, lackenby2020algorithms3manifoldtheory] for lying exactly midway between the well-understood surfaces and the wildly behaved 4-manifolds. For example, several 3-manifold problems were related to important questions in complexity theory: $S^{3}$ recognition was proved by Schleimer to be in NP [schleimer2011sphere], whereas approximations of some quantum invariants are firmly tied to quantum computing and are known to be $\texttt{P}-$hard [alagic2011quantum, alagic2014quantum]. Similarly, the problems of classifying [costantino2024learning, kuperberg2018identifying] and manipulating the geometry and topology of these spaces [burton2013new, lackenby2024triangulation] have also received significant theoretical and practical attention, with implementations of different methods publicly available in the most important software of the field, regina [regina] and SnapPy [SnapPy].
The algorithmic point of view is particularly relevant for the construction of censuses of homomorphism classes of 3-manifolds [burton2004efficient, burton2014cusped, Burton2020_TabulationKnots3Manifolds, burton2015edge]. Drawing inspiration from the related problem of knot classification (this time, up to ambient isotopies), where the cumulative work of researchers [burton2020next, conway1970enumeration, dowker1983classification] has culminated in the widely accessible census of $\sim 350^{\prime}000^{\prime}000$ prime knot diagrams, computational methods are expected to foster the elaboration of tables of 3-manifolds, particularly valuable assets in the proposition and verification conjectures, as well as in the discovery of counterexamples.
Abstracting modern approaches [burke_et_al:LIPIcs.SoCG.2025.28, burton2014cusped, burton2020next, burton2015edge] of the construction of censuses of topological spaces, one requires 1. compact data structures to represent massive amounts of (candidate) spaces and the iterative construction of all spaces up to a maximum complexity; 2. efficient operations to disqualify inputs (such as composite knots, links, non-prime 3-manifolds) and simplify the representation combinatorics (Reidemeister moves, Pachner moves, etc); and, 3. methods for discriminating spaces, such as invariants. For 3-manifolds, these steps are often addressed through triangulations—homeomorphic copies described by a finite set of tetrahedra and gluing rules among their faces—of complexity usually measured by the number of tetrahedra. On the other hand, theoretical bounds on the minimal number of tetrahedra to represent some spaces [datta2007minimal, jaco20030, jaco2006layered, matveev2007algorithmic] and the difficulty of generating particular structures, such as hyperbolicity [matveev2007algorithmic], can limit the effectiveness of triangulation-based censuses.
@@
Heegaard splittings emerge as an alternative to triangulations. They depict a closed 3-manifold $M$ as the result of gluing two better-behaved pieces called handlebodies along a common boundary. Splittings can also be combinatorially described by Heegaard diagrams, two sets of curves in a surface whose intersection pattern conveys much of the topological information about $M$. We suggest here a new data structure to represent Heegaard diagrams based on normal coordinates, of a size as large as the binary of the number of intersections between the two sets of curves. The gain in measuring the complexity of 3-manifolds through this notion is not to be overlooked: we present in Section 5 families of manifolds whose diagrams’ complexity grows linearly with the size of the input, but whose triangulations necessarily use exponentially many tetrahedra. However, what is more important is that due to the proliferation of efficient operations in normal coordinates [Erickson2013, lackenby2024some, schaefer2002algorithms, schaefer2008computing, Stefankovic], we can establish polynomial-time algorithms even in this compressed notion of complexity, leading to basic manipulations of Heegaard splittings and the computation of simple 3-manifold invariants such as the fundamental and homology groups. Moreover, as we point out in Remark 4.11, some important complexity results about 3-manifolds—such as sphere (co-)recognition—are still preserved if we assume compressed inputs, which tells us much about the inherent difficulty of these problems. We hope that these techniques are useful for the study and classification of topological spaces that are beyond the reach of efficiency when represented by other methods.
@@
The paper begins with a background section on both low-dimensional and computational topology (Section 2), where we review some of the concepts and results that will be needed in the later parts. We refer the reader to [schultens2014introduction] and [singer2015lecture] for basic ideas of geometric topology, such as homeomorphisms, isotopies, gluing maps, and manifolds. Our contribution starts in Section 3, where we formally present our data structure for Heegaard splittings. Section 4 establishes basic algorithms in this structure. We compare early implementations [^1] The code is publicly available as the FHDPy package at https://github.com/HLovisiEnnes/FHDpy. Details of the implementation are also outlined in Appendix B. of some of these algorithms with triangulation-dependent methods in Section 5. Finally, Section 6 points to open questions towards the construction of censuses of 3-manifolds. We also present appendices with extensions of the techniques described in the main text and details on the implementation.
@@
Before we close this introduction, we draw the reader’s attention to two subtleties of the current work. First, although most of our tools apply to generalized Heegaard splittings defined for orientable compact 3-manifolds, potentially with boundary [johnson2006notes, schultens2014introduction], we will focus solely on closed 3-manifolds. Therefore, all 3-manifolds here will be assumed orientable and closed, unless explicitly stated otherwise. Second, we assume throughout a unit-cost RAM model with bit-size $W$ [hagerup1998sorting]. This means that the four integer operations and integer comparison can be executed in constant time, provided that the numbers involved have binary representations of size at most $W$.
@@ 2.1. Background on low-dimensional topology
By surface, we understand a compact and connected smooth 2-manifold $\Sigma$ with or without boundary. All surfaces will be assumed orientable and oriented. If $\Sigma$ is closed, it is uniquely given, up to homeomorphism, by an integer $g$ called the surface’s genus [farb2011primer] and we denote this surface by $\Sigma_{g}$. A (simple closed) curve and an arc are images of proper embeddings of $S^{1}$ and $[0,1]$ in $\Sigma$, respectively. A multicurve is a finite collection of disjoint curves. Given a multicurve $\gamma$, we denote its number of connected components by $\gamma$ and, with some abuse of notation, write $c\in\gamma$ if $c$ is a component of $\gamma$. Whenever possible, we reserve Greek letters for multicurves and Latin letters for curves and arcs.
We will most often consider the isotopy class of a (multi)curve $\gamma$ in $\Sigma$, $\gamma]$. A class $\gamma]$ is called essential if no component is isotopic to a point. From now on, we will not differentiate $\gamma$ from its isotopy class $\gamma]$ in the notation if there is no confusion.
Figure 1. The recursive relation that defines the Lickorish generators of the mapping class group.
Let $\gamma$ and $\delta$ be two oriented (multi)curves in $\Sigma$, isotoped to intersect transversely. For each intersection point, we let its index be +1+1 if the orientation of the curves agrees with the orientation of $\Sigma$ as in the right-hand rule, otherwise we set it to 1-1. The algebraic intersection number of $\gamma$ and $\delta$, $\hat{i}(\gamma,\delta)$, is the sum of the indices of all their transverse intersections; the algebraic intersection number is well defined up to isotopy [farb2011primer]. Two (multi)curves are said to be in efficient position if they are representative of their respective isotopy classes such that the total number of intersection points between them is minimal. Equivalently, they are in efficient position if they do not have segments that cobound disks, known as bigons, on the surface [farb2011primer].
The mapping class group of the surface $\Sigma_{g}$, $\operatorname{Mod}(\Sigma_{g})$, is the group of orientation-preserving $\Sigma_{g}\rightarrow\Sigma_{g}$ homeomorphisms up to isotopy. For each $g$, $\operatorname{Mod}(\Sigma_{g})$ is finitely generated as a group by an alphabet $\mathcal{L}$, consisting of Dehn twists about 3g13g-1 (isotopy classes of) canonically defined essential curves $s_{1},\dots,s_{3g-1}$ [lickorish1964finite] (Figure 1). Formally, a Dehn twist $\tau_{s}:\Sigma_{g}\rightarrow\Sigma_{g}$ about $s$ is given by cutting out a cylindrical regular neighborhood around the curve, giving a $2\pi$ left twist to one of the cylinder’s boundaries and gluing the cylinder back into the surface. We call the curves $s_{1},\dots,s_{3g-1}$ the surface’s Lickorish generators. Hopefully without confusion, we also call the Dehn twists in the alphabet $\mathcal{L}=\{\tau_{s_{1}},\dots,\tau_{s_{3g-1}}\}$ Lickorish generators.
@@
$\mathcal{H}_{\gamma}=\Sigma_{g}\times[0,1]\cup_{\gamma\times\{0\}}2\text{-handles}\cup 3\text{-handles}$
where we attach 2-handles along the curves $\gamma$ in $\Sigma\times\{0\}$ and fill any spherical boundary component with 3-handles. Note that $\partial\mathcal{H}_{\gamma}=\Sigma_{g}$ and that the handlebody $\mathcal{H}_{\gamma}$ is homeomorphic to the connected sum of $g$ solid tori. The components of $\gamma$ are meridians of $\mathcal{H}_{\gamma}$ in the sense that they bound disks in the handlebody’s interior. Two systems $\gamma$ and $\gamma^{\prime}$ in the same surface $\Sigma_{g}$ will be called equivalent if $\gamma^{\prime}$ is a set of meridians in $\mathcal{H}_{\gamma}$ (which implies that $\gamma$ is a set of meridians in $\mathcal{H}_{\gamma^{\prime}}$).
@@
In this paper, we restrict the analysis to minimal systems $\gamma$, that is, we assume $\gamma=g$. In recent decades, it became common to extend the definition of a system to any multicurve in the surface that defines a handlebody [hempel20013, yoshizawa2014high]. In fact, the data structure described in Section 3 can be generalized to this notion of system, and there are algorithms to translate from and to the special case of cardinality $g$. For completeness, we describe these extensions in the Appendix A.
Let $c_{1},c_{2}$ be two fixed components of a system $\gamma$ and $p$ an arc from a puncture corresponding to $c_{1}$ to a puncture corresponding to $c_{2}$ in $\Sigma_{g}\backslash\gamma$. We define the band sum of $c_{1},c_{2}$ and $p$ as the curve $c$ in $\Sigma_{g}\backslash\gamma$ indicated by the surgery in Figure 2. One can show that $c$ is essential and bounds a disk in $\mathcal{H}_{\gamma}$ [ennes2025hardnesscomputationquantuminvariants, johnson2006notes], implying that $\gamma^{\prime}=\gamma-\{c_{1}\}\cup\{c\}$ is a system equivalent to $\gamma$. We call this whole operation a disk slide. With finitely many applications of disk slides and isotopies, a system $\beta$ can be transformed into any equivalent system $\beta^{\prime}$ [johnson2006notes].
@@
A genus $g$ Heegaard splitting is the closed 3-manifold resulting from the gluing of two handlebodies with boundaries homeomorphic to $\Sigma_{g}$ through a gluing map $f:\Sigma_{g}\rightarrow\Sigma_{g}$. In fact, one might assume that $f$ is given by a $\phi\in\operatorname{Mod}(\Sigma_{g})$ composed with some canonical orientation-reversing homeomorphism. This means that a Heegaard splitting can be given by a pair $(\Sigma_{g},\phi)$ where $\phi$ is a word on Lickorish generators. We call this presentation of Heegaard splitting a Heegaard word. Every closed 3-manifold has a Heegaard splitting and can then be associated with some pair $(\Sigma_{g},\phi)$ [moise1952affine, saveliev2011lectures].
@@
Although naturally combinatorial, a Heegaard word $(\Sigma_{g},\phi)$ hardly conveys topological information about the closed 3-manifold $M$ it represents. Nonetheless, if we fix a system of meridians $\alpha$ on one of the two handlebodies, $\mathcal{H}_{\alpha}$, then $\beta=\phi(\alpha)$ is a system of meridians of the other handlebody, $\mathcal{H}_{\beta}$. We call the tuple $(\Sigma_{g},\alpha,\beta)$ a Heegaard diagram. We note that Heegaard diagrams are algorithmically independent of Heegaard words: any two systems $\alpha$ and $\beta$ will define the same splitting, provided that the curves are meridians in their respective handlebodies. Diagrams are therefore not unique for a given splitting. For example, if $\beta$ and $\beta^{\prime}$ are isotopic diagrams on $\Sigma_{g}$—that is, $\beta$ and $\beta^{\prime}$ are isotopic multicurves in $\Sigma_{g}$—then $(\Sigma_{g},\alpha,\beta)$ and $(\Sigma_{g},\alpha,\beta^{\prime})$ naturally define the same manifold. Similarly, if $\beta$ is equivalent to $\beta^{\prime}$, we say that the diagrams are equivalent and note that $(\Sigma_{g},\alpha,\beta)$ and $(\Sigma_{g},\alpha,\beta^{\prime})$ also define the same 3-manifold.
@@
Stabilizations provide yet another method to generate alternative Heegaard splittings of the same 3-manifold. Formally, let $(\Sigma_{g_{1}},\alpha_{1},\beta_{1})$ be a Heegaard diagram of a 3-manifold $M_{1}$ and $(\Sigma_{g_{2}},\alpha_{2},\beta_{2})$ be a diagram of $M_{2}$. The diagram $(\Sigma_{g_{1}+g_{2}},\alpha_{1}\cup\alpha_{2},\beta_{1}\cup\beta_{2})$ represents the connected sum $M_{1}\#M_{2}$, where $\Sigma_{g_{1}+g_{2}}$ is the connected sum of $\Sigma_{g_{1}}$ and $\Sigma_{g_{2}}$ taken along open disks disjoint from $\alpha_{1}$$\alpha_{2}$, $\beta_{1}$, and $\beta_{2}$. In particular, noting that $S^{3}$ has a splitting with diagram $(\Sigma_{1}=T^{2},m,\ell)$, where $|m\cap\ell|=1$, and that $M\#S^{3}$ is homemorphic to $M$, the splitting $(\Sigma_{g+1},\beta\cup\{\ell\})$ represents the same manifold as $(\Sigma_{g},\alpha,\beta)$. We say that the former is a stabilization of the latter. In the other direction, a Heegaard diagram $(\Sigma_{g},\alpha,\beta)$ represents a stabilization of a lower genus splitting if and only if there is a meridian $a$ in $\mathcal{H}_{\alpha}$ and a meridian $b$ in $\mathcal{H}_{\beta}$ such that $|a\cap b|=1$ [johnson2006notes]. In particular, we say that this pair of meridians is a stabilization pair and that the diagram is stabilized. For a Heegaard diagram $(\Sigma_{g},\alpha,\beta)$, $a$ and $b$ form a trivial stabilization pair if $a$ is a component $\alpha$, $b$ is a component of $\beta$, they form a stabilization pair, and $b\cap a^{\prime}=\emptyset$ and $b^{\prime}\cap a=\emptyset$ for any other component $a^{\prime}\in\alpha$ and $b^{\prime}\in\beta$. Whenever $a$ and $b$ form a trivial stabilization, cutting the surface along $a$ induces a Heegaard diagram in the surface of genus one unit smaller. This process will be called destabilization.
@@
A splitting $(\Sigma_{g},\mathcal{H}_{\alpha},\mathcal{H}_{\beta})$ is called reducible if there is a curve $c$ in $\Sigma_{g}$ that is a meridian of both $\mathcal{H}_{\alpha}$ and $\mathcal{H}_{\beta}$. We call a diagram $(\Sigma_{g},\alpha,\beta)$ reducible if either (1) there are a component $a\in\alpha$ and a component $b\in\beta$ that are isotopic in $\Sigma_{g}$; or (2) if there is an essential separating curve $s$, disjoint from both $\alpha$ and $\beta$, and such that $\s\Sigma_{g}\backslash$ s defines a connected sum of two Heegaard splittings. The uniqueness of fundamental group decomposition [aschenbrenner20153] and [schultens2014introduction][Theorem 6.3.5] implies that a reducible diagram represents a reducible splitting. A 3-manifold is reducible if it contains a sphere that does not bound a ball. If $(\Sigma_{g},\alpha,\beta)$ is stabilized, then either the splitting is reducible or a stabilization of a genus 1 splitting of $S^{3}$; moreover, a reducible splitting is stabilized or a splitting of a reducible manifold [johnson2006notes].
@@
Useful algebraic information on the 3-manifold can also be readily obtained from its Heegaard diagrams. Label the curves $a_{1},\dots,a_{g}\in\alpha$ and $b_{1},\dots,b_{g}\in\beta$ and give each of them an orientation. It follows from an application of the Van Kampen theorem [hatcher2002algebraic, hempel20223, johnson2006notes] that we can compute a presentation $\langle x_{1},\dots,x_{g}|r_{1},\dots,r_{g}\rangle$ for the fundamental group $\pi_{1}(M)$ by traversing the curve $b_{i}$ along the assumed orientation and, whenever it intersects some $a_{j}$, we append $x_{j}$ to the relation $r_{i}$ if the index of the intersection is positive, otherwise we append $x_{j}^{-1}$ to it (if $b_{i}$ does not intersect $\alpha$, $r_{i}$ is just the trivial relation).
@@
Because the first (integral) homology group of $M$, $H_{1}(M)$, is the abelianization of $\pi_{1}(M)$ [hatcher2002algebraic], we can present $H_{1}(M)$ as $\langle x_{1},\dots,x_{g}|r^{\prime}_{1},\dots,r^{\prime}_{g},[x_{i},x_{j}]\rangle$, where each $r^{\prime}_{i}=x_{1}^{k_{i}^{1}}\dots x_{g}^{k_{i}^{g}}$ can be computed by commuting the generators $x_{j}$ in the relations $r_{i}$. Note that $k_{i}^{j}=\hat{i}(b_{i},a_{j})$ and we store this information in a $g\times g$ presentation matrix for $H_{1}(M)$, $K=[k_{i}^{j}]$, whose Smith normal form carries all the necessary information about this Abelian group. For example, the (first) Betti number of $M$ is given by the number of zero-rows. A 3-manifold with Betti number equal to zero is called a rational homology sphere, and it will be an integral homology sphere if the Smith form of the presentation matrix is the identity.